to win a lotto in one state, one must correctly select 6 numbers from a collection of 46 numbers. the order does not matter. How many different selections are possible
Answer by Maths68(1474)

If order does not matter then we can use following formula to find different COMBINATIONS of 6 numbers out of 46 numbers
Use following Combination formula
nCr = n! / r!(n-r)!
n=46
r=6
=46!/6!(46-6)!
=46!/[6!(40)!]
=(46*45*44*43*42*41*40!)/(6*5*4*3*2*1)(40!)
Cancel out 40!
=46*45*44*43*42*41/(6*5*4*3*2*1)
=6744109680/720
=9366819
Possible different selection 9366819